You wish to test the following claim (HaHa) at a significance level of α=0.10α=0.10.

      Ho:μ1=μ2Ho:μ1=μ2
      Ha:μ1<μ2Ha:μ1<μ2

You believe both populations are normally distributed, but you do not know the standard deviations for either. However, you also have no reason to believe the variances of the two populations are not equal. You obtain the following two samples of data.

Sample #1 Sample #2
82.9 76 98.2
63.9 76.2 86.9
71.7 82.5 77.4
87.4 61.8 85.1
89
83 88.5 86.4
78.9 92.7 88.2
92.7 71.6 88
89.5 96.4 87.1
80.7 86.6 74.3
84.4 80.7 80.5
86.4 81.2 73.8
93



What is the test statistic for this sample? (Report answer accurate to three decimal places.)
test statistic =

What is the p-value for this sample? For this calculation, use the degrees of freedom reported from the technology you are using. (Report answer accurate to four decimal places.)
p-value =
Answer #1

Let x Sample 1 and y. Sample 2 S.No. ly-ybar)2 3.0786 14.0280 2.7073 34.2763 63.1294 4 SE54u (x-xbar)2 829 8.8619 76 15.390713 12 = 104.3119 i For sample 2 , M=13, Ex= 1039 F = $x = 1034 – = 79.9231 S? E (x-70)2 - 1251-7431_ 15,2= 104,319 For samplesp=562,-1) + S₂2 (M2-1) nitt-2 =104.3119x12 + 43. 1902×21 13+22-2 = 2158.318 33 sp= 65.4036 op vri, this JB tán = 0.3498 Sp V